# Robustly optimal dynamics for active matter reservoir computing

### v1

@@ III.3 Emergent mechanisms in the optimal regime
Here, we visually investigate the information processing mechanisms behind RC in the optimal (nearly critically damped) regime. For the optimal regime under Lorenz-63 driving, two alternating dynamical states are consistently observed: In Fig. 4, we observe a quasi-stationary response state. The system forms a moving interface synchronous with the slow driver, enabled largely by the overdamped dynamics at short time scales, yet still showing (subtle) long-wavelength fluctuations in structure and velocities. Upon rapid changes in the driver’s dynamics, which occur around the transitions between the two attractors, synchronization breaks and a multi-step feedback cycle are triggered. This cycle amplifies signals of the rupture, converts them into different visible mechanisms, and eventually restores the quasi-stationary, driver-synchronized state. Note that we used a larger system ($N=500$) for visual purposes, which showed very similar phenomenology, both qualitatively and quantitatively, to the smaller default system ($N=200$), even though the local density is slightly higher.
Quasi-stationary response: When the driver moves slowly (in Fig. 4, upper panel), an interface with marked surface tension is formed. This interface creates an exclusion zone, a region of ultra-low-density gas, around the driver. The interface moves synchronously with the driver and remains stable as long as the driver moves at slow enough speeds. This ability of the system to adapt to a driver even as the latter moves at non-negligible speeds may be enhanced by the active propulsion. Notably, long-range wave motion can be seen in the videos colored by speed (Supplementary Video Tab. 2/1), or in the videos of velocity fluctuations (Supplementary Video Tab. 4/2). We hypothesize that the localization of the driver position through agents with high speeds around the interface could constitute beneficial swarm dynamics for the high reservoir computing performance. Likewise, the long-range oscillatory motion of the swarm induced by the driver could mean a consistent information transfer, which would be useful for reservoir computing.
Highly nonlinear response: When there is abrupt driver motion (Fig. 4, lower panel), the driver first touches the interface, whereby agents experience a strong local repulsion force. The whole system experiences a sudden perturbation, showing oscillatory behavior. As the interface breaks, the driver moves faster than agents may coherently respond, causing local shear thinning. An aisle briefly forms through the bulk of the agents, followed by viscoelastic self-healing. Once the driver slows down at a spatially separated position, the local shear thinning caused by the driver in the bulk allows for the efficient formation of the next quasi-stationary interface. The end of this cycle effectively erases the memory effects of the most recent rupture.
@@ III.4 Correlative spatiotemporal velocity fluctuations indicate high performance regimes
A different viewpoint is considering the strength of anti-correlations at the first minimum (at $r_{min}$) of the CVC in Fig. 5b. We observe that at the optimal parameter combination (triangle symbol), the greatest degree of anti-correlative strength is reached. It offers a complementary indicator for high predictive performance compared to our calculated dynamical susceptibility, particularly at the crossover region between underdamped/ overdamped and optimally damped regimes (the yellow dashed diagonals), respectively. At the upper-right-hand corner of the diagram, we observe extrema in $\chi$ and $\text{CVC}(r_{\text{min}})$, around the “L”/hexagon symbol. Strong rotational modes (see Fig. 15b) that occur in these regimes and that were not accounted for when computing velocity fluctuations could explain these extrema.
@@
Investigating the velocity fluctuation snapshots in Fig. 5c shows that in the underdamped regime (cross symbol), the amount of correlation among velocity fluctuations is low. As the speed-controlling force increases (diamond symbol), the velocity fluctuations become more correlated, indicated by a smoother color gradient of the arrows – a more gradual change of orientation of neighboring velocity fluctuations. Only near the optimal and overdamped cases, represented by the teal triangle and green circle symbols, do we visually recognize the localization of prominent anti-correlated velocity fluctuations. Here, exclusively agents close to the source of the disturbance, the driver, experience strong velocity fluctuations. The remainder of the swarm features velocity fluctuations on a much smaller scale and forms a neutral background, allowing a sharp distinction of velocity fluctuations induced by the driver. In contrast, in the underdamped regime (cross symbol), driver-induced velocity fluctuations cannot be distinguished from intrinsic velocity fluctuations of the system. This highlights the importance of the following duality for active matter reservoir computing: a well-defined swarm “ground state” and local, driver-induced agent excitations. This finding solidifies Lymburn et al.’s observation that there must be a functional dependency of the active matter system on the input signal [63] – which is here stronger in the critically damped regime compared to the underdamped regime.
@@
In general, anti-correlated velocity fluctuations could form via the propagation of a (longitudinal) wavefront from a central source of disturbance, from an elastic-type backlash or rebound, or from a backflow towards a common center from different angles. The CVC calculates correlations at the same time-points just like the dynamic susceptibilities do. Thus, the anti-correlative behavior arising from longitudinal wavefronts around the driver’s source is probably the strongest contributor to the full signal, as seen also in the visuals of the velocity fluctuations (Fig. 5c). Anti-correlative velocities that show elastic recoil or rebound effects with time-delay could contribute to the VAC negative-valued peaks in Fig. 3c; incoherent contributions coming from backflow or backlash of momenta manifesting at diverse delayed timescales should broaden or distort peaked signals in the VAC Fig. 3c.
@@ III.5 Robust optimality across various classes of chaotic driving
To investigate this inquiry, we empirically choose four additional time-series benchmarks belonging to different dynamic classes [31]: the chaotic attractors Hénon-Heiles [87] (twists), Rössler [88] (rotations), Chua [89] (lobes), and Lorenz-96 [90] (ripples). While these attractors belong to different dynamical classes and feature different Lyapunov exponents, we normalize their trajectories such that the drivers’ mean speeds match that of the previously studied Lorenz attractor (see Appendix A for details). The Lyapunov times of the chaotic trajectories indicate the varying difficulties of the prediction problems. Hence, to compare RC predictive performance across chaotic attractors, we align the predicted time ahead to the Lyapunov time predicted ahead for the Lorenz-63 system studied earlier ($\approx 0.45283\ t_{\text{lyap}}^{\text{L63}}$) and shown in Fig. 2.
@@
At the circle symbols, the Rössler attractor can be predicted with a performance of up to 0.9862, Chua with up to 0.9828, Hénon-Heiles with up to 0.9560, and Lorenz-96 with up to 0.8737. When comparing the underdamped and critically damped regimes, we observe marked differences for attractors of varying difficulty. The underdamped regime still yields useful reservoir dynamics for less “difficult” chaotic attractors (Rössler, Hénon–Heiles, Chua), compared to the critically damped regime, with performance differences of around 0.10. For the chaotic attractors Lorenz-63 and Lorenz-96, we observe higher differences of 0.30 between the two regimes. While the numerical difference between 0.90 and 0.990 might seem small, there is a significant difference in the smoothness of the prediction quality, as seen in Fig. 27. The predicted, less challenging time series of Rössler, Hénon–Heiles and Chua are resolved by our near-critically damped active matter reservoirs much better compared to their underdamped counterparts. The amplitudes of the fluctuations in the predicted time series are much smaller.
@@
The surprisingly high prediction scores for the underdamped regimes for less chaotic attractors could have another reason beyond the lower prediction difficulty. Each chaotic attractor naturally drives the soft matter system into out-of-steady-state configurations distinct from those of the Lorenz-63 (and -96) systems, which may be more or less favorable for making predictions. For example, for the Rössler attractor, we measure high scalar rotation parameters in the underdamped regime (see Fig. 22b. This is because the Rössler driving protocol induces strong milling in the underdamped swarm (see Supplementary Video Tab. 10/4). These sustained rotations increase the structural and dynamical order of the swarm, which could improve predictive performances via a higher consistency of swarm responses in this particular driving context. Similar considerations could hold for Chua (underdamped oscillations, see Supplementary Video Tab. 10/6) and possibly Hénon-Heiles (underdamped pulsations, see Supplementary Video Tab. 10/2). This observation highlights the intricacies and variability of non-equilibrium dynamics in the underdamped regime, which vary more strongly for each prediction problem and depend on the driving context. Yet, even the critically damped regime displays non-equilibrium phenomenology different from that of the Lorenz-63 case: the distinction between quasi-stationary and highly non-linear response (Fig. 4) is less visible in the corresponding videos. (Note, however, that we here display videos for slightly higher underdamping (circle symbol)). Moreover, the time-scales of particular emergent mechanisms like shear thinning are not transient, but occur over prolonged time-scales with a weaker degree of intensity.
@@
We also provide predictive performances for a constant predicted real-time ahead (instead of constant Lyapunov time ahead) of $\Delta t_{\text{pred}}=0.50$ for all attractors in Fig. 26.
@@
A high dynamical susceptibility, in the way it was calculated, is still a good indicator for the critically damped regime (see Fig. 24). By accounting for the swarm rotations and dilatations as proposed in Ref. [94], a properly corrected dynamical susceptibility could yield a more generic indicator for the high-performance region for benchmarks from different dynamical classes. Refer to Figs. 21 - 25 for an overview of observables for the four studied benchmark systems.
@@ III.7 Revisiting the role of particle alignment
Higher interaction radii $r_{a}$ and alignment strengths $K_{a}$ then lead to the formation of an active crystal droplet phase (triangle symbol). This phase yields performances of $P=0.798$ and shows high polar order (see Fig. 29a). From Supplementary Video Tab. 12/1 we observe that the long-ranging weak alignment force creates a condensed swarm. Here, the alignment force effectively suppresses driver-induced agent inertia, and repelled agents quickly return to the flock. This leads to the formation of an effectively critically damped droplet with more consistent dynamics upon driving, which is a desirable property for reservoir computing with soft matter as analyzed in Sect. III.1. One reason why the performance could be lower here compared with the actual critically damped regime (shown in Fig. 2a) is that the droplet is not confined to the center of the simulation box. Due to the high mean agent speed, a result of the chosen speed controller settings, the droplet rotates around the box center as an active crystal. This could reduce the prediction quality because similar driving patterns could lead to different swarm states due to this inherent rotational swarm motion. Furthermore, we note an interesting property of the flock: for slow driving, we observe a rotating active crystal that avoids the driver, but is only weakly perturbed by it. For abrupt driving, the active crystal flock aligns its direction of motion to that of the driver. This additional information processing mechanism could improve the predictive quality of the droplet because information about recent driving is conserved in the swarm.
Further increasing radius and strength then leads to alignment force contributions that dominate over the other force terms. This creates a locally frustrated swarm motion defying the homing force (square symbol). The frustrated swarm splits when crossing the (periodic) box boundaries. We hypothesize that the waning influence of the driver due to frustrated motion, hence creating different patterns for similar driving histories, reduces the predictive performance in this regime. For extremely high alignment strengths and radii, we observe a globally frustrated swarm state (nabla symbol) with performances close to zero. The freezing of the initial orientation is indicated by an absence of order, measured by polarity values close to zero (Fig. 29a). We note that for this speed-controller setting, the dynamical susceptibility has its maximum and the connected velocity correlation function has its first local minimum in the locally frustrated regime (square symbol), not in the optimal droplet regime. The mean squared displacement at the Lorenz-63 Lyapunov time is low in the optimal regime but features another minimum at the transition from locally to globally frustrated regimes. This hints that these indicators are not absolute and must be applied and interpreted with care.
@@
We conclude that for a swarm with high mean speed (enforced by Lymburn et al. (2021) speed-controller settings), weak and long-ranged alignment forms an active crystal droplet which may improve predictive performance. Nevertheless, the high predictive performance in the critically damped regime in the absence of any alignment interaction suggests that even passive matter could be a suitable candidate for soft matter reservoir computing.
@@ III.8 The optimal regime is robust against variations in homing force strength
Decreasing the homing force strength significantly in the critically damped regime (represented by the diamond symbol) effectively eliminates the agent’s pull toward the center of the simulation box. Because of this, agents distribute themselves across the simulation box. The agent-agent repulsive forces cause them to occupy the entire simulation box while organizing themselves into a highly ordered crystalline structure. The repulsive force of the driver cuts out a vacancy area in the crystal. This area is not immediately reoccupied by agents due to the relative weakness of agent-agent repulsions. Strong and abrupt driving causes traveling waves through the bulk, which can be nicely observed in the corresponding video Supplementary Video Tab. 13/4. This phononic information transfer is yet another example of the rich phenomenology found in our driven soft matter systems. The relatively low performance of $P=0.642$ at this parameter combination may be due to the lack of a strong interface that ensures continued information transfer from the driver to the system under slow driving. Small, quasi-stationary driving changes the system state only minimally. On top of that, the overall heterogeneity of (collective) agent responses becomes limited compared to the critically damped droplet, which should lower the responsiveness and thus expressivity of the soft matter system for representing and processing relevant information. A complementary parameter scan over homing force strength $K_{h}$ and target agent speed $s$ with a fixed speed-controller strength is shown in Fig. 30.

### v2

@@ III.3 Emergent mechanisms in the optimal regime
Here, we visually investigate the information processing mechanisms behind RC in the optimal (nearly critically damped) regime. For the optimal regime under Lorenz-63 driving, two alternating dynamical states are consistently observed: In Fig. 4, we observe a quasi-stationary response state. The system forms a moving interface synchronous with the slow driver, enabled largely by the overdamped dynamics at short time scales, and marginally showing subtle long-wavelength oscillations in structure and velocities. Upon rapid changes in the driver’s dynamics, which occur around the transitions between the two attractors, synchronization breaks and a multi-step feedback cycle are triggered. This cycle amplifies signals of the rupture, converts them into different visible mechanisms, and eventually restores the quasi-stationary, driver-synchronized state. Note that we used a larger system ($N=500$) for visual purposes, which showed very similar phenomenology qualitatively and quantitatively to the smaller default system ($N=200$), even though the local density is slightly higher.
Quasi-stationary response: When the driver moves slowly (in Fig. 4, upper panel), an interface with marked surface tension is formed. This interface creates an exclusion zone (a vacuum) around the driver. The interface moves synchronously with the driver and remains stable as long as the driver moves at slow enough speeds. This ability of the system to synchronize an interface to a driver even as the latter moves at intermediate speeds might be enabled by the active propulsion. Notably, long-range wave motion can be seen in the videos colored by speed (Supplementary Video Tab. 2/1), or in the videos of velocity fluctuations (Supplementary Video Tab. 4/2). We hypothesize that the localization of the driver position with an interface could constitute a main beneficial collective mechanism for the high reservoir computing performance.
Highly nonlinear response: When there is abrupt driver motion (Fig. 4, lower panel), the driver first touches the interface, whereby agents experience a strong local repulsion force. The whole system experiences a sudden perturbation, showing oscillatory behavior. As the interface breaks, the driver moves faster than agents may coherently respond, causing local shear thinning. A corridor briefly forms through the bulk of the agents, followed by self-healing of the old interface via viscous flow. Once the driver slows down at a spatially separated position, the local shear thinning caused by the driver in the bulk allows for the efficient formation of the next quasi-stationary interface. The end of this cycle effectively erases the memory effects of the most recent rupture.
@@ III.4 Correlative spatiotemporal velocity fluctuations indicate high performance regimes
A different viewpoint is considering the strength of anti-correlations at the first minimum (at $r_{min}$) of the CVC in Fig. 5b. We observe that at the optimal parameter combination (triangle symbol), the greatest degree of anti-correlative strength is reached. It offers a complementary indicator for high predictive performance compared to our calculated dynamical susceptibility, particularly at the crossover region between underdamped/ overdamped and optimally damped regimes (the yellow dashed diagonals), respectively. At the upper-right-hand corner of the diagram, we observe extrema in $\chi$ and $\text{CVC}(r_{\text{min}})$, around the “L”/hexagon symbol. Strong rotational modes (see Supplementary Fig. 15b) that occur in these regimes and that were not accounted for when computing velocity fluctuations could explain these extrema.
@@
Examining the velocity fluctuation snapshots in Fig. 5c shows that in the underdamped regime (cross symbol), the amount of correlation among velocity fluctuations – in their spatial distribution – is low. As the speed-controlling force increases (diamond symbol), the velocity fluctuations become more correlated, indicated by a smoother color gradient of the arrows – a more gradual change of orientation of neighboring velocity fluctuations. Only near the optimal and overdamped cases, represented by the triangle and circle symbols, do we visually recognize the localization of prominent anti-correlated velocity fluctuations. Here, those agents that are close to the source of the disturbance, the driver, experience strong velocity fluctuations. The remainder of the swarm features velocity fluctuations on a much smaller scale, creating a neutral background for a sharp contrast of velocity fluctuations localized near the driver. In contrast, in the underdamped regime (cross symbol), driver-induced velocity fluctuations cannot be distinguished as easily from intrinsic velocity fluctuations of the system. This highlights the favorability of the following duality for active matter reservoir computing: a well-defined swarm “ground state” and local, driver-induced agent excitations. This finding solidifies Lymburn et al.’s observation that there must be a functional dependency of the active matter response on the input signal [62] – which is here stronger in the critically damped regime compared to the underdamped regime.
@@
In general, correlated velocity fluctuations could form in several ways, such as the propagation of a wavefront or pulse. These are tied to basic viscous and elastic, or mechanical properties of the fluid. The CVC calculates correlations at the same time-points just like the dynamic susceptibilities do. The anti-correlative behavior arising from wavefronts or pulses forming around the driver’s source is probably the strongest contributor to the full signal, as seen also in the visuals of the velocity fluctuations (Fig. 5c). In the absolute VAC’s negative-valued peaks in Fig. 3c, incoherent contributions manifesting at diverse timescales can broaden or distort the shape Fig. 3c. Depending on the types of interactions allowed, these correlations can also arise differently, though [97].
@@ III.5 Robust optimality across various classes of chaotic driving
To investigate this inquiry, we empirically choose four additional time-series benchmarks belonging to different dynamic classes [31]: the chaotic attractors Hénon-Heiles [89] (twists), Rössler [90] (rotations), Chua [91] (lobes), and Lorenz-96 [92] (ripples). While these attractors belong to different dynamical classes and feature different Lyapunov exponents, we normalize their trajectories such that the drivers’ mean speeds match that of the previously studied Lorenz attractor (see Appendix A for details). The Lyapunov times of the chaotic trajectories indicate the varying difficulties of the prediction problems. Hence, to compare RC predictive performance across chaotic attractors, we predict the same Lyapunov time ahead as for the Lorenz-63 system studied earlier ($\approx 0.45283\ t_{\text{lyap}}$, different real times) and shown in Fig. 2.
@@
At the circle symbols, the Rössler attractor can be predicted with a performance of up to 0.9862, Chua with up to 0.9828, Hénon-Heiles with up to 0.9560, and Lorenz-96 with up to 0.8737. When comparing the underdamped and critically damped regimes, we observe marked differences for attractors of varying difficulty. The underdamped regime still yields quite useful reservoir dynamics for less “difficult” chaotic attractors (Rössler, Hénon–Heiles, Chua). Compared to the critically damped regime, the performance differences are around 0.10. For the chaotic attractors Lorenz-63 and Lorenz-96, we observe higher differences of 0.30 between the two regimes. While the numerical difference between 0.90 and 0.990 might seem small, there is a significant difference in the smoothness of the prediction quality, as seen in the Supplementary Fig. 27. The predicted, less challenging time series of Rössler, Hénon–Heiles and Chua are resolved by our near-critically damped active matter reservoirs much better compared to their underdamped counterparts. The amplitudes of the fluctuations in the predicted time series are much smaller.
@@
The surprisingly high prediction scores for the underdamped regimes for less chaotic attractors could have another reason beyond the lower prediction difficulty. Each chaotic attractor naturally drives the soft matter system into out-of-steady-state configurations distinct from those of the Lorenz-63 (and -96) systems, which may be more or less favorable for making predictions. For example, for the Rössler attractor, we measure high scalar rotation parameters in the underdamped regime (see Supplementary Fig. 22b. This is because the Rössler driving protocol induces strong milling in the underdamped swarm (see Supplementary Video Tab. 10/4). These sustained rotations increase the structural and dynamical order of the swarm, which could improve predictive performances via a higher consistency of swarm responses in this particular driving context. Similar considerations could hold for Chua (underdamped oscillations, see Supplementary Video Tab. 10/6) and possibly Hénon-Heiles (underdamped pulsations, see Supplementary Video Tab. 10/2). This observation highlights the intricacies and variability of non-equilibrium dynamics in the underdamped regime, which vary more strongly for each prediction problem and depend on the driving context. Yet, even the critically damped regime displays non-equilibrium phenomenology different from that of the Lorenz-63 case: the distinction between quasi-stationary and highly non-linear response (Fig. 4) is less visible in the corresponding videos. (Note, however, that we here display videos for slightly higher underdamping (circle symbol)). Moreover, the time-scales of particular emergent mechanisms like shear thinning are not transient, but occur over prolonged time-scales with a weaker degree of intensity.
@@
We also provide predictive performances for a constant predicted real-time ahead (instead of constant Lyapunov time ahead) of $\Delta t_{\text{pred}}=0.50$ for all attractors in Supplementary Fig. 26.
@@
A high dynamical susceptibility, in the way it was calculated, is still a reasonable indicator of the vicinity of the critically damped regime (see Supplementary Fig. 24). By accounting for the swarm rotations and dilatations as proposed in Ref. [96], a properly corrected dynamical susceptibility could yield a more generic indicator for the high-performance region for benchmarks from different dynamical classes. Refer to Supplementary Figs. 21 - 25 for an overview of observables for the four studied benchmark systems.
@@ III.7 Revisiting the role of particle alignment
Higher interaction radii $r_{a}$ and alignment strengths $K_{a}$ then lead to the formation of an active crystal droplet phase (triangle symbol). This phase yields performances of $P=0.798$ and shows high polar order (see Supplementary Fig. 29a). From Supplementary Video Tab. 12/1 we observe that the long-ranging weak alignment force creates a condensed swarm. Here, the alignment force effectively suppresses driver-induced agent inertia, and repelled agents quickly return to the flock. This leads to the formation of an effectively critically damped droplet with more consistent dynamics upon driving, which is a desirable property for reservoir computing with soft matter as analyzed in Sect. III.1. One reason why the performance could be lower here compared with the actual critically damped regime (shown in Fig. 2a) is that the droplet is not confined to the center of the simulation box. Due to the high mean agent speed, a result of the chosen speed controller settings, the droplet rotates around the box center as an active crystal. This could reduce the prediction quality because similar driving patterns could lead to different swarm states due to this inherent rotational swarm motion. Furthermore, we note an interesting property of the flock: for slow driving, we observe a rotating active crystal that avoids the driver, but is only weakly perturbed by it. For abrupt driving, the active crystal flock aligns its direction of motion to that of the driver. This additional information processing mechanism could improve the predictive quality of the droplet because information about recent driving is conserved in the swarm.
Further increasing radius and strength then leads to alignment force contributions that dominate over the other force terms. This creates a locally frustrated swarm motion defying the homing force (square symbol). The frustrated swarm splits when crossing the (periodic) box boundaries. We hypothesize that the waning influence of the driver due to frustrated motion, hence creating different patterns for similar driving histories, reduces the predictive performance in this regime. For extremely high alignment strengths and radii, we observe a globally frustrated swarm state (nabla symbol) with performances close to zero. The freezing of the initial orientation is indicated by an absence of order, measured by polarity values close to zero (Supplementary Fig. 29a). We note that for this speed-controller setting, the dynamical susceptibility has its maximum and the connected velocity correlation function has its first local minimum in the locally frustrated regime (square symbol), not in the optimal droplet regime. The mean squared displacement at the Lorenz-63 Lyapunov time is low in the optimal regime but features another minimum at the transition from locally to globally frustrated regimes. This hints that these indicators are not absolute and must be applied and interpreted with care.
@@
We conclude that for a swarm with high mean speed (enforced by Lymburn et al. (2021) speed-controller settings), weak and long-ranged alignment forms an active crystal droplet which improves predictive performance. Nevertheless, the high predictive performance in the critically damped regime in the absence of any alignment interaction suggests that even passive matter could be a suitable candidate for soft matter reservoir computing.
@@ III.8 The optimal regime is robust against variations in homing force strength
Decreasing the homing force strength significantly in the critically damped regime (represented by the diamond symbol) effectively eliminates the agent’s pull toward the center of the simulation box. Because of this, agents distribute themselves across the simulation box. The agent-agent repulsive forces cause them to occupy the entire simulation box while organizing themselves into a highly ordered crystalline structure. The repulsive force of the driver cuts out a vacancy area in the crystal. This area is not immediately reoccupied by agents due to the relative weakness of agent-agent repulsions. Strong and abrupt driving causes traveling waves through the bulk, which can be nicely observed in the corresponding video Supplementary Video Tab. 13/4. This phononic information transfer is yet another example of the rich phenomenology found in our driven soft matter systems. The relatively low performance of $P=0.642$ at this parameter combination may be due to the lack of a strong interface that ensures continued information transfer from the driver to the system under slow driving. Small, quasi-stationary driving changes the system state only minimally. On top of that, the overall heterogeneity of (collective) agent responses becomes limited compared to the critically damped droplet, which should lower the responsiveness and thus expressivity of the soft matter system for representing and processing relevant information. A complementary parameter scan over homing force strength $K_{h}$ and target agent speed $s$ with a fixed speed-controller strength is shown in Supplementary Fig. 30.